### v3

# PATCH: a deep learning method to assess heterogeneity of artistic practice in historical paintings

PATCH: a deep learning method to assess heterogeneity of artistic practice in historical paintings
In the Renaissance and Early Modern period, paintings were largely produced by master painters who directed workshops of apprentices and others who often contributed to the piece. Discerning who created these masterworks and how they did so is a central question in technical art history1 and a nontrivial problem that machine learning can help solve by extending analysis to a microscopic scale2. Analysis of workshop paintings presents a challenge, however, because information about the members of workshops and the processes by which artworks were created remains elusive3; 4. Thus, external examples are not available to train networks to recognize. Here we present a novel machine learning approach we call pairwise assignment training for classifying heterogeneity (PATCH) that is capable of identifying individual artistic practice regimes with no external training data. We apply this method to two historical paintings by the Spanish Renaissance master, El Greco, and our findings regarding one of the works potentially challenge previous studies that assert that a considerable portion of the painting was completed by workshop members after El Greco’s death5. PATCH outperforms statistical and unsupervised machine learning methods in this complex pairwise comparison problem lacking “ground truth” data, making it potentially useful across similar cases in the social and natural sciences, including image segmentation in remote sensing6, urban development and design7, and anomaly detection manufacturing contexts8; 9, among others.
There are many ways to create a painting. The process can be collaborative or solitary, and there is no end to the styles, tools, materials and techniques that can be implemented. The evolution of what we call artistic practice—the technical, stylistic, material, and even physiological aspects of an artist’s creative process—over time and its variation between artists remains a key element of technical art history scholarship. From the late Middle Ages through the Early Modern period, on the other hand, painting was typically done in a bustling workshop10. A master painter would establish a workshop and assemble a group of apprentices, journeymen and specialist painters and even collaborate with associates to help fulfill commissions and create smaller works to support the enterprise between larger contracts.
Workshop practice adds a second layer of complexity to artistic practice. Master painters varied considerably in their managerial styles. Some performed much of the work themselves, while others dictated the composition but left many elements to the hands of workshop members (e.g., van Dyck11). Still others involved apprentices in the conceptualization of a piece12. Often, there was a hierarchy, with young apprentices typically assigned more mundane tasks such as the mixing of paints and the construction of canvases and more senior workshop members directed to paint portions of larger works in the master’s style3.
Because workshops varied so widely in their organization and management yet produced works with cohesive styles, deeper understanding of variation in artistic practice can shed light on workshop practice and vice versa. Research into these key aspects of Renaissance and Early Modern art production is hindered by a lack of surviving textual sources regarding the structure, size, and membership of individual workshops as well as their day-to-day operation and the variation in artistic practices3. While scholars have been able to reconstruct individual artists’ workshops through a variety of methods, our collective understanding of the inner workings of workshops is still fragmentary4. Artistic practice varies widely, encompassing materials, methods, styles, and individual contributions. Understanding who interacted with individual paintings and how they did so—discerning the artists at work in a painting, the interplay of their individual “hands,” and the interaction of the artist’s hand and eye with paint and brush13—is an important open question in art history1.
Machine learning has recently emerged as a promising complement to traditional art historical analyses in applications including preservation, conservation, and the detection of forgeries14; 15; 16; 17, as well as in fighting illegal trafficking in antiquities18. Much work involving style is focused on broad classification of works19; 20, such as the use of a neural network to arrange works of the Western canon in the correct chronological order21. Recently published work by some of the authors of the present study demonstrated the efficacy of ML for attribution of paintings by analyzing high-resolution topographic images of the surface texture of oil paintings by known authors (student painters) in a controlled experimental setting. In a supervised machine learning analysis, a convolutional neural network (CNN) was trained to sort microscopic surface textures among artists, assigning 1 cm2 patches of paintings to the correct authors with 95% accuracy2. Thus, the method has shown that microscopic analysis can complement traditional art historical analyses of macroscopic features.
Unfortunately, supervised deep learning on topographic images cannot be used when there is no ground truth for the network to learn on. When the object of analysis is a painting produced by a Renaissance or Early Modern artist’s workshop, the number of artistic practice regimes is generally unknown, known examples of the included artists’ work may not exist, and knowledge of the painting practices employed is incomplete. Unsupervised learning, wherein the machine analyzes individual objects (e.g., images or patches of an image) and creates classes based on the statistical properties of those objects, would appear more appropriate. However, a naive unsupervised analysis still depends on hyperparameters that can be tuned to yield larger or smaller numbers of classes. Whether performed by the trained art historian or by a neural network, the problem of identifying regions of shared artistic practice is, in a word, nontrivial.
Here, we demonstrate a novel technique, to our knowledge, for ML-based image attribution, and use it to assign paintings to their respective artist(s). We then combine this method with network analysis to identify artistic practice regimes (combinations of artists and materials) and create a measure of the heterogeneity of artistic practice (HAP) within a given painting or set of paintings. We apply this method to two important historical paintings by El Greco, one considered to be entirely by the master himself, and one previously thought to feature the work of members of his workshop.
The method described here, Pairwise Assignment Training for Classifying Heterogeneity (PATCH), uses supervised learning toward an unsupervised end. Rather than train a network to recognize individual classes, we test whether the network is capable of learning to distinguish between two objects and then use that information to build post-hoc classes. The PATCH method relies on the inability of the network to correctly sort patches of paintings or regions of a painting painted by the same artist. Given patches of two paintings by different painters—a different-artist pair—the network will learn to assign those patches to the correct painting with a high degree of accuracy, as demonstrated in our previous publication. But if the network is given patches of two paintings by the same artist—a same-artist pair—it will not be able to learn to assign those patches to the correct painting; it will perform no better than a coin flip. We then create a network of same-artist pairs and use network analysis to construct classes corresponding to artistic practice regimes. In this way, we obviate the need for separate ground truth information for individual artists and artistic practice regimes, enabling successful applications in contexts such as workshop paintings where the ground truth is not known with certainty.
Development and validation of the PATCH algorithm
The PATCH method (Figure 1) consists of two phases. The first phase, pairwise assignment training, identifies pairs of paintings or pairs of regions of a single painting that are the “same,” that is, created by the same artist under the same conditions. The second phase, community finding, creates post hoc classes based on these pairings.
The first phase was developed on a set of paintings by nine known student artists. Each artist individually painted three paintings, all using the same tools, materials, and subject (a photograph of lilies), as described previously2. While the artists were instructed to realistically depict the subject, choices of style and technique were left up to the artists. One artist was only able to produce two paintings, and one painting was removed from the sample due to the optical profilometry data being corrupted, for a total of 25 paintings. Topographical information for each painting was recorded using high resolution spectral confocal optical profilometry with a spatial resolution of 50 microns, and a height repeatability of 200 nm. The resulting data were processed to remove any large-scale warping of the canvas by subtracting a mean-filtered version of the height map (with filter radius of 0.5 cm)2. Because in this dataset the paintings were small (12 cm 15 cm), each “region” chosen for the pairwise training consists of an entire painting. The region is then subdivided into 1 cm2 square patches, which corresponds to the size where the earlier supervised machine learning analysis 2 was able to achieve maximum classification accuracy. This patch size is big enough to contain significant information about brushstrokes, while yielding sufficiently large datasets for training (in this case 180 patches per painting).
The network architecture chosen was VGG-16, implemented in TensorFlow and pretrained on the ImageNet dataset. Given the computational complexity of the training—the combinatorics of many region-to-region comparisons, repeated over many folds, described in detail below—the relatively light-weight but accurate VGG-16 architecture was ideal for the PATCH approach (see the Supplementary Materials for more information on hyperparameter settings and optimization). In order to measure the ability of the network to learn to distinguish between two paintings or regions of a painting, we ran 26 training folds. Each fold comprised the following steps: 1 cm2 patches from each painting/region were randomly selected, with replacement. Sampling with replacement creates a bootstrapping effect that drives down the success rate of same-artist pairs and drives up the success rate for different-artist pairs. To eliminate the influence of directional elements, the corners of each patch were trimmed to create an octagon, and the patch was randomly rotated to one of eight possible orientations. The network was then trained for 25 epochs with validation set size set to 30% of patches, and the maximum validation accuracy (percentage of validation patches correctly identified) was recorded. The distribution of the 26 maximum validation accuracies was then compared to the distribution expected if the network were assigning patches at random.
If the network has failed to learn to assign objects to their correct classes, it should perform this task no better than if it were assigning them randomly. We are measuring the maximum validation accuracy from 25 training epochs in each of 26 folds. Thus, the distribution expected if the network fails to learn to distinguish between painters would be the distribution , the probability that you will see a maximum of m heads (i.e., correct assignments) over k repetitions (k epochs) of an experiment where you flip a fair coin n times (the number of patches in the test set), which we call the random assignment distribution (RAD) (see the Supplementary Materials for an analytical derivation of this distribution).
Two decision criteria were selected to determine whether the distribution of maximum validation accuracies indicated that the network had failed to learn to distinguish between the input paintings: 1) if the observed mean had a z-score less than 2 relative to the RAD mean and 2) if the largest of the observed maximum validation accuracies (the right edge of the observed distribution) among the 26 folds was less than an empirically determined threshold based on finite sampling from RAD, plus 10% to account for the effects of bootstrapping (selection with replacement, see Supplementary Information for more details). Using these decision criteria, the classification performance in determining whether the artist was the same or different was exceptional. We identified same-artist pairs (n = 23) and different-artist pairs (n = 277) with F1 scores of .889 and .991 respectively (see Table S2). The PATCH algorithm substantially outperformed a statistical method based on surface roughness, and beat a variety of unsupervised and supervised clustering approaches (see details in SI Appendix C) in F1 for both same-artist and different-artist pairs.
While pairwise assignment training performs exceptionally well in identifying artists with control for materials, situations where ground truth about the artistic practices employed is not available require some means to create the post hoc classes to which regions of a painting will be assigned. To this end, we use network analysis. We construct a network where same-artist pairs (or, for historical paintings, same-practice pairs) of paintings or regions are connected by an edge (lines connecting regions in Figure 3) and different-artist/practice pairs are unconnected.
The PATCH method is concerned with “sameness,” meaning we need a way to remove spurious edges from the network. To accomplish this, we employ an edge pruning process based on uniqueness. We assume that regions with a high degree (number of connections) have more common features and nodes with a low degree have more unique features. Therefore, a connection between regions with more unique features is more likely to be correct. We operationalize this assumption by assigning each edge a score based on the average uniqueness of its nodes, where uniqueness is defined as the percentage of nodes a node does not connect with: where max = the maximum possible degree (i.e., number of connections) and degree = the degree of one of the two nodes in question. There were two erroneous same-artist pairs in our experimental dataset (9% of edges in the network), and these edges had the lowest uniqueness scores. Thus, we trim the 9% of edges (or more if there are ties) in a given network with the lowest uniqueness scores.
With potentially spurious edges removed and the remaining edges weighted equally, we then employ a community finding algorithm to identify groups of regions that are the most similar. If a set of regions has a large number of internal links within the set and fewer links to outside regions, it forms a “community” within the network22; 23; 24, suggesting that the regions in a given community were likely painted under different circumstances (artists, materials, etc.) than those in other communities. To establish the degree of difference between communities, we can characterize the community structure of the network via measures such as modularity ()—the fraction of edges in the network that are internal to communities minus the mean fraction in a network with the same communities but where the edges are completely randomized23. We implemented the Louvain community finding algorithm25 in Gephi26. With the resolution set to 1.0, the algorithm returns the partition of the network with the maximum modularity. The Louvain algorithm returned a correct partition of the experimental dataset (nine disjoint communities for nine different artists) with (see Figure S4).
El Greco, The Baptism, and Christ on the Cross
El Greco, born Domenikos Theotokópoulos (1541-1614), is regarded as a pillar of the Renaissance in Spain and as an early progenitor of modernism27. El Greco became a master of icon painting in Crete before journeying through Venice and Rome, eventually settling in Toledo, Spain5; 28. It was there that the master married Byzantine and Venetian motifs29 and developed the peculiar style for which he is most widely recognized, with its distorted human figures, “expressive hands”30, and ”exploitation” of ”pure colors to their limits”31. Two examples of that style are The Baptism of Christ (1624, Hospital Tavera, Toledo, Esp., henceforth, The Baptism), shown in Figure 2A, and Christ on the Cross with Landscape (ca. 1600-1610, Cleveland Museum of Art, Cleveland, OH, henceforth Christ on the Cross), shown in Figure 2B.
These works were chosen for this analysis because of the difference in how art historians have characterized their authorship. The Baptism has long been thought to feature the work of El Greco and at least one other artist. Primary historical evidence indicates that El Greco began the painting under a contract with the Hospital Tavera but retained it at his death in 1614. It was delivered to the Hospital nearly a decade later32, during which period art historians have proposed the painting was finished by workshop members, particularly the master’s son, Jorge Manuel5. Previous art historical studies have attempted to attribute regions of The Baptism to the master and others (illustrated in Figure S3) largely through connoisseurship—visual analysis of artists’ styles and artistic choices33. El Greco himself is proposed to have painted the entire top, with the possible exception of the robe of the angel on the right (which was not scanned for this study due to its poor condition)32; 34, as well as the angel in green (bottom left), with the exception of the wings. Jorge Manuel is proposed to have painted John the Baptist (bottom right) and the adjacent figure in red. Lopera34 and Wethey32 ascribe the image of Christ (bottom center) to Jorge Manuel and El Greco, respectively, suggesting that it could represent the work of both artists. Lopera believes that a third hand may have been involved in rendering the faces of the angels in the background on the bottom. The landscapes at the bottom were extended and “transformed into a river” by an unknown artist sometime between receipt of the painting and its installation in the epistle-side altarpiece ca. 1660, which was not an unusual practice at the time. Restoration of the landscape on the right at the very bottom was performed in 193632; 34; 35. In contrast, Christ on the Cross has been entirely attributed to El Greco himself. It is generally accepted that El Greco created many works in their entirety without the aid of his workshop, including Christ on the Cross. Francis36 states that most of El Greco’s approximately twenty renditions of the Crucifixion were “by his own hand,” and notes the “lightning dexterity of brushwork” in this particular example.
Application to El Greco’s The Baptism
Two large sections of The Baptism were scanned in situ using the Factum Foundation’s Lucida scanner. The Lucida scanner is a non-contact laser triangulation scanner with 100-micron lateral resolution and is capable of acquiring scans in situ over a large area by stitching together 48 cm square tiles. The scanned regions used in this study are indicated in Figure 3. The height data was processed in the same way as for the student paintings described above, including correction for possible canvas warping.
Regions of the topographic image of The Baptism were selected by hand in order to minimize variation in subject matter (e.g., faces, robes) within each region and to constrain region size to between 180 and 540 cm2. The lower bound was chosen because the method had been developed and validated using 180-patch student paintings and the upper bound was selected to ensure adequate representation of patches when regions of significantly different size were compared. We avoided selecting areas with significant cracking or damage, including the robes of the large angels to the left and right of God, and the red swatch held by the “green angel” in the bottom half of the painting. When comparing regions of different sizes during the PATCH analysis, the sample size was set equal to the number of patches in the smaller region.
Pairwise comparisons of all 55 regions (a total of 1485 pairwise tests) of The Baptism yielded a network with 356 edges, which was reduced to 314 edges after pruning. The Louvain community-finding algorithm25 with resolution set to 1.0, returned a partition with that featured four communities (Figure 3a). A value of is often considered evidence of structure in real-world networks23.
The four communities are mapped onto the painting in Figure 3B. The first community overlaps much of the top portion of the painting, including the image of God the Father. The second overlaps the image of John the Baptist and Christ’s torso and hands. The third covers much of the center of the painting, the face of Christ and the faces of many of the background angels. The final community is confined to Jesus’ legs and the rocks at the bottom of the painting.
Application to El Greco’s Christ on the Cross
As in the case of The Baptism, regions of the topographic image of Christ on the Cross were selected by hand in order to minimize variation in subject matter and to constrain region size. Areas with apparent cracking or damage, such as the background and trees on the lower left and the cross beam to the immediate left of Christ’s face, were not included.
Pairwise comparison of all 24 regions (276 pairwise tests) yielded a network with 123 edges, which was reduced to 108 edges after pruning. The Louvain community finding-algorithm25 at resolution 1.0 generated a partition with that featured two communities. Two is the smallest possible number of communities for a non-fully connected graph, as compares inter- and intra-community connections. Communities are mapped onto the painting in Figure 4. Constituent regions of both communities are not randomly distributed, with one community overlapping much of Christ’s torso and face, while the other community overlaps Christ’s legs, the lower portion of the cross and regions of the background.
Discussion
To our knowledge, ours is the first study to demonstrate the PATCH learning method and apply it both to paintings from a controlled experiment and to historical paintings by a well-known artist with a workshop. Among AI/ML-based methods designed to aid in the analysis of historical paintings, PATCH is notable in that it requires no external, ground truth data. Networks trained on known examples can accomplish tasks associated with workshop practice and authorship with exceptional accuracy. For example, Ugail and colleagues recently used transfer learning to authenticate paintings by the Renaissance master Raphael (Raffello Sanzio)37. By combining edge detection with typical feature extraction in a residual neural network, they were able to recognize works by Raphael with 98% accuracy. However, known examples of artists’ work may not be available, as exemplified by the workshop context. Our method shows exceptional promise as a tool for ML analysis of complex historical works where known samples of work by the artists involved do not exist. Indeed, it may have extensive application in other image analysis tasks where there is little to no ground truth information available.
Further, the second phase of PATCH creates what we consider a measure of the heterogeneity of artistic practice (HAP) within a given dataset. Communities found within a dataset represent different artists, materials, or both. The degree to which those communities are interconnected provides information about the possibility of shared authorship, materials, or other practices. A fully connected network with no distinct communities has . As the number of communities increases and the interconnection between communities decreases, modularity increases toward 1 (a completely disconnected network has undefined modularity, but would effectively represent the maximum heterogeneity). The networks from our three studies provide empirical examples of lower and higher modularity networks. Christ on the Cross and The Baptism have values that cluster around 0.3, considered the threshold for evidence of structure, indicating low heterogeneity. The network for the experimental student paintings has much greater heterogeneity at (see Figure S4). The maximum modularity depends on the number of edges in the network38, but also on the number of communities in the network. A network with two fully connected and completely disjoint communities has , for example. As such, network modularity can function as a measure of absolute observed heterogeneity (Figure 5).
HAP informs interpretation of our analyses of Christ on the Cross and The Baptism. For the former, our analysis identifies two communities of regions, but with below the threshold generally accepted to indicate structure (0.231 vs 0.3). The regions in each community do cluster in space, however, which is worth considering. One community overlaps Christ’s head and torso and surrounding areas. Conservation files generously provided by the Cleveland Museum of Art indicate that the figure of Christ (areas overlapped by the community in yellow in Figure 4) contains inpainting from previous conservation, which could create a distinguishable signal, though damage to the canvas is another, more likely culprit. Further mapping of damage and past conservation efforts will improve our understanding of this work. However, the results comport with the notion of Christ on the Cross as the work of a single artist with some variation introduced by artistic practice or possibly early conservation practice.
At first glance, the communities identified in our analysis of The Baptism appear to roughly correspond to the proposed attributions by previous authors discussed above. The community in yellow in Figure 3 overlaps much of the top of the painting, previously attributed to El Greco, while the community in blue overlaps John the Baptist and the figure in red, assigned to Jorge Manuel. We also find a combination of communities in the face and torso of Christ, which was attributed to both artists. A single, small and particularly unique community (in orange in Figure 3) overlaps the landscape at the bottom of the painting and Christ’s legs.
Overall, however, the PATCH findings suggest a different story underlying the existing art historical interpretation. The partition of The Baptism has , which is only slightly above the threshold for evidence of structure. Nearly a third of all edges in the network are interconnections between three of the four communities we have identified. The yellow and green groups have 49 intercommunity edges, blue and green 30, and blue and yellow 22. If the blue, green, and yellow communities indeed represented three distinct artists working with the same or similar materials, we would expect the communities to be more insular and the modularity of the network to be substantially higher, similar to what we observed in our controlled student painting dataset. Yet the modularity of The Baptism is much closer to that of Christ on the Cross (), which appears to feature one artist. It is worth noting that, as shown in the analysis of student paintings, the CNN is particularly proficient at recognizing different-artist pairs and is more likely to incorrectly label a same-artist pair as different than the converse. Thus, the intercommunity edges in The Baptism network that survived our initial pruning are not liable to be spurious.
This suggests that some unifying factor connects these three communities. One possibility is a single artist working with different brushes. Renaissance artists used brushes made from the bristles or hair of several different mammal species, variation in the characteristics (e.g., coarseness) of which could affect the deposition of paint 39; 40; 41 (see Supplementary Materials for more information). The evidence could also indicate the work of one artist with changing style or technique over time. Interestingly, Biannuci and colleagues42 present evidence that El Greco suffered a series of ischemic events, one in the 1590s and a second in 1608 (the year the contract for the altarpieces for the Hospital Tavera was initiated), “resulting in progressive disabilities with fluctuating course characterized by temporary improvements and worsening before his death.” The regions that were previously identified as the work of different individual painters could conceivably represent variation in the master’s individual style (however this may have arisen) over the course of his final years. Investigation of changes in style and technique and their effect on PATCH analysis will be an important direction for future research.
PATCH has the capability to make a substantial contribution to research as a complement to existing art historical methods. The accuracy of the method has been demonstrated in the validation on student paintings. Our analyses of Christ on the Cross and The Baptism show the true potential of PATCH by contributing important new information to art historical scholarship on the creation of objects in El Greco’s workshop in Toledo. In the analysis of The Baptism, the fact that the orange group, which overlaps a region known to have been altered after the delivery of the painting, is so disconnected from the remainder of the communities (see Supplementary Materials for further analyses) speaks to the method’s capability of recognizing a unique contribution or substantial departure from the technique used in the rest of the painting. That the other three communities roughly correspond to art historical attributions suggests PATCH possesses a kind of semantic fidelity43—our method recognizes areas that a trained art historian would classify as different. However, the connections between those areas revealed by PATCH also call those attributions into question. Computer vision allows us to view the surface of a painting at a different scale. Previous analyses indicate that features as small as the diameter of a single paint brush bristle (length scales of 0.2–0.4 mm) may be integral to the network’s identification of an artist2. By analyzing spatial correlations at this microscopic scale, the network may be revealing aspects of the physics of paint application and of the physiology of hand movements. PATCH adds micro-scale features to the macro (brushstrokes) and meta-scale (historical and material) data, allowing for a full-scale analysis and the discovery of heretofore unseen evidence of the processes by which paintings were created.
Future research with PATCH will focus on teasing apart the influences of materials and artists on the heterogeneity of artistic practice. Expansion of PATCH applications to include intra-painting comparisons and other art historical applications will be important as well. However, uses of PATCH are not limited to those contexts: it provides a general-purpose approach to assessing regional similarity in images or data that can be encoded in image-like arrays. As such, PATCH is a promising method for applications where supervised learning is impossible (because the ”ground truth” is unknown) and alternative unsupervised approaches are ineffective, especially for a posteriori cluster number. Potential applications span the physical and social sciences, and include image segmentation tasks in applications such as medical imaging44, agricultural remote sensing6, urban development and design7, and microstructural analysis, as well as anomaly detection in a range of manufacturing contexts8; 9, among others.
Data availability
The dataset used to develop and validate the PATCH method is available on OSF at osf.io/a5xnh/. Height data from The Baptism and Christ on the Cross are the property of the institutions that own those paintings: the Fundación Casa Ducal de Medinaceli and the Cleveland Museum of Art, respectively.
Code availability
The code used to run the PATCH algorithm is available on GitHub: https://github.com/hincz-lab/PATCH.
Acknowledgments
The authors would like to thank Fundación Casa Ducal de Medinaceli for providing data for The Baptism and Sr. Juan Manuel Albendea Solis for his feedback on our initial draft, which helped us reevaluate our conclusions and improve the manuscript. We would also like to thank the Cleveland Museum of Art for use of the data from Christ on the Cross, and the Lapis Senior Conservator of Paintings and Head of Paintings Conservation, Dean Yoder, for his help and feedback. The authors acknowledge support from the Expanding Horizons Initiative of Case Western Reserve University (CWRU) College of Arts and Sciences and the Jack, Joseph and Morton Mandel Foundation. Computational support was provided by the CWRU High Performance Computing cluster and a workstation donated through the Dell Seed Program.
Author contributions
AV conceived the PATCH method; LNS and MSM conceived the initial project; AV, MH, KS, MSM developed the method; AI oversaw creation of paintings; ALo, CB, and ITM performed scans; AV and MM wrote PATCH code; AV, ALi, MM and MH performed experiments and analyses; AV, MH, KS, ESB, CP, ALi, and LNS wrote the manuscript.
Ethics declarations
Competing interests: The authors declare no competing interests.
Supplementary information
Supplementary Information is available for this paper.
Correspondence
Correspondence and requests for materials should be addressed to AV.
Supplemental Materials:
“PATCH: a deep learning method to assess heterogeneity of artistic practice in historical paintings”
Appendix A: Analytical derivation of the random assignment distribution
In order to determine whether the network has learned to correctly sort patches of both images, we generated the distribution of correct assignments expected if the network were assigning each patch to a class at random (as though it were simply flipping a coin). The network is trained for 26 folds of 25 epochs each. In each epoch, the network is trained on 70% of the patches and tested on the remaining 30% (there are equal numbers of patches from each image). The percentage of this test set that is assigned correctly is called the “validation accuracy.” At the end of each fold, the maximum validation accuracy is recorded, the dataset is resampled, and a new fold begins. After all folds are completed, we have a distribution of 26 maximum validation accuracies that represent how well the network has learned to distinguish between the two images in each of its attempts. The derivation of this expected distribution is as follows:
The goal is to find the distribution , the probability that you will see a maximum of heads over repetitions (“ epochs”) of an experiment where you flip a fair coin times. We start with the distribution of finding heads in a single experiment, which is the binomial distribution for a fair coin,
where is the binomial coefficient. We also define a tail distribution, which is the cumulative probability that you would find or more coin flips in a single experiment:
Here is a hypergeometric function, defined by the series:
where for and . There are other equivalent ways of expressing the tail probability of a binomial distribution (using for example incomplete beta functions), but all of them are similarly complicated.
If we repeat the experiment times, the probability that every single experiment had fewer than heads is:
The difference between consecutive values of this probability, , is the probability that we would find a maximum exactly equal to over repetitions. Hence we get our final answer:
It is easy to check that this probability is properly normalized:
In practice, we compared the mean and maximum of the observed distribution to the random assignment distribution derived above. However, the theoretical maximum of the distribution is m = n. Given enough tries (infinite sampling), a random assignment engine would eventually achieve perfect accuracy. Thus, we developed an empirical “maximum” value based on finite sampling using the image size of our student paintings (180 patches, 360 total, n = 108 patches in the test set). We created a simple coin-flipping script that gathered the maximum “validation accuracies” (i.e., percentage of flips that land heads) for 25x108 flips and ran 250,000 iterations of this process. This generated a maximum value of m = 80 “correct assignments,” or 74.074%. The probability of 80 correct assignments is 0.000004676581 (truncated for clarity). As the distribution is a probability mass function based on discrete events, for tests with larger numbers of patches, we used the number of correct assignments with a probability closest to this probability, even if it was less probable.
Appendix B: Hyperparameter Settings and Optimization
The first phase of PATCH analysis, pairwise assignment training, was performed using VGG16 architecture pretrained on the Imagenet database. Output from VGG-16 was pooled and flattened and passed through two dense layers with dropout rates of 0.25. The learning rate was set to 0.0001 and the batch size was set to 32 patches.
The model hyperparameters for the neural network artistic practice discriminator model are optimized with a commercially available Bayesian hyperparameter optimization package built around the tree parzen optimization algorithm 45. The optimization loss is based on the model performance after initial training on a restricted subset of the training data, because using the full dataset for this initial optimization step is computationally infeasible. In particular each model was trained for: 6 folds, 10 epochs, and patch size 200 on a reduced student painting dataset consisting of 5 paintings with 4 same artist combinations and 6 different artist combinations. Training loss was calculated as the RMSE distance between the predicted artist similarity and the ground truth, averaged over all painting combinations. The restricted training scheme was employed to reduce the total optimization time on finite computational resources. The considered hyperparameters and ranges are shown in Table S1. The model structure consists of a variable number of dense, fully-connected layers following the VGG16 model, where the number of nodes per layer decreases linearly from a variable maximum at the first dense layer to 2 at the output layer. Maximum posterior estimation was performed on the generated set of all trained models (130 instances) to predict the optimal hyperparameter set. Final optimization of the model shown in the manuscript utilized a full training scheme on the full dataset, as described in the manuscript, and is performed by hand. This is necessary since the full training scheme leads to unreasonably long convergence time in the optimization algorithm. The final hyperparameters align closely with the maximum posterior estimation result, as shown in Figure S1. Note that the largest deviation between the optimized and maximum posterior hyperparameters occurs for the learning rate, which is expected since the longer number of epochs in the full training process allows the model to converge with a smaller learning rate.
| Hyperparameter | Min search val. | Max search val. | Max posterior | Final optimization |
| Learning rate | 0.0001 | 0.02 | 0.0025 | 0.0001 |
| Dense layers | 1 | 4 | 1 | 1 |
| Max nodes | 25 | 200 | 97 | 64 |
| Dropout rate | 0.0 | 0.25 | 0.20 | 0.25 |
| L2 reg. rate | 0.0001 | 0.13 | 0.0007 | 0.001 |
Appendix C: PATCH in Comparison to Alternative Methods
The PATCH algorithm outperformed alternative methods (described below) for the student painting dataset, identifying same-artist pairs (n = 23) and different-artist pairs (n = 277) with excellent classification performance, summarized in Table S2. In the table, precision and recall are calculated by treating each category (same-artist pairs or different-artist pairs) as “positives” in turn. For example, for the same-artist pairs, the true positives TP number of same-artist pairs classified as same, false positives FP number of different-artist pairs classified as same, and false negatives FN number of same-artist pairs classified as different. The precision = , the recall = , and the F1 score is the harmonic mean of precision and recall. The definitions for the different-artist pair category are as above, but with “same” and “different” interchanged. The last three columns of the table show averages of the same-artist and different-artist columns for precision, recall, and F1. The PATCH method was superior to the tested alternatives in the majority of metrics, notably achieving the highest F1 scores for same-artist pairs and different-artist pairs. The closest competitor, supervised UMAP with a Fourier feature extractor, achieved only marginally higher precision for same-artist pairs and recall for different-artist pairs, beating PATCH by 2.9% and 0.3% respectively. However it performed substantially worse on same-artist recall (0.652 versus 0.87), thus scoring lower on F1. That method’s marginally increased precision is therefore outweighed by its inability to identify same-artist pairs (15/23, 6̃5%). In the following we describe the alternative methods we tested.
| Same-artist pairs () | Different-artist pairs () | Average | |||||||
| Method | Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 |
| PATCH | 0.909 | 0.87 | 0.889 | 0.989 | 0.993 | 0.991 | 0.949 | 0.931 | 0.940 |
| Surface roughness | 0.250 | 0.826 | 0.384 | 0.982 | 0.794 | 0.878 | 0.616 | 0.810 | 0.631 |
| Fourier / UMAP | 0.257 | 0.783 | 0.387 | 0.978 | 0.812 | 0.888 | 0.618 | 0.797 | 0.637 |
| Fourier / sUMAP | 0.8 | 0.522 | 0.632 | 0.961 | 0.989 | 0.975 | 0.881 | 0.755 | 0.803 |
| Fourier / UMAP | 0.254 | 0.652 | 0.366 | 0.967 | 0.841 | 0.9 | 0.611 | 0.747 | 0.633 |
| Fourier / sUMAP | 0.938 | 0.652 | 0.769 | 0.972 | 0.996 | 0.984 | 0.955 | 0.824 | 0.877 |
| Fourier / UMAP | 0.274 | 0.739 | 0.4 | 0.975 | 0.838 | 0.901 | 0.624 | 0.788 | 0.65 |
| Fourier / sUMAP | 0.688 | 0.478 | 0.564 | 0.958 | 0.982 | 0.97 | 0.823 | 0.73 | 0.767 |
| CNN / UMAP | 0.197 | 0.565 | 0.292 | 0.957 | 0.809 | 0.877 | 0.577 | 0.687 | 0.584 |
| CNN / sUMAP | 0.37 | 0.739 | 0.493 | 0.976 | 0.895 | 0.934 | 0.673 | 0.817 | 0.713 |
| CNN / UMAP | 0.23 | 0.609 | 0.333 | 0.962 | 0.83 | 0.891 | 0.596 | 0.72 | 0.612 |
| CNN / sUMAP | 0.34 | 0.739 | 0.466 | 0.976 | 0.881 | 0.926 | 0.658 | 0.81 | 0.696 |
| CNN / UMAP | 0.239 | 0.696 | 0.356 | 0.97 | 0.816 | 0.886 | 0.604 | 0.756 | 0.621 |
| CNN / sUMAP | 0.348 | 0.696 | 0.464 | 0.972 | 0.892 | 0.93 | 0.66 | 0.794 | 0.697 |
Surface roughness method
Pairwise comparison of large numbers of regions via PATCH is computationally expensive, so a simpler method would be preferable if it were equally accurate. One simple way of distinguishing between regions of a painting using topography would be to compare the “roughness” of the surfaces. We applied a roughness-based method to the student paintings to determine its accuracy relative to the PATCH method. We used the standard deviation of height values as a measure of roughness, and calculated the standard deviation of height values for each 1x1 cm2 patch of each student painting. We then compared paintings pairwise using the pairwise Wilcoxon rank sum test implemented in R using the stats package. A significant result on the pairwise Wilcoxon test means the distributions (specifically the medians) are significantly different, which in this context would indicate a different-artist pair. Thus, same-artist pairs would be tests where the null hypothesis could not be rejected. Using the Bonferroni correction for multiple tests with =0.01, any pairwise comparison with a p-value greater than 0.00003 was considered a same-artist pair. This method resulted in F1 scores of 0.384 for same-artist pairs and 0.878 for different-artist pairs, substantially worse than PATCH. This also indicates that PATCH is not simply sorting patches of images based on their roughness, but using more complex characteristics of the height data for classification.
Unsupervised UMAP methods
As described in the main text, PATCH achieves unsupervised ends by supervised means. How would PATCH fair against fully unsupervised learning on the surface topography, using clustering of dimension-reduced feature vectors for each patch? To facilitate this comparison, we developed an unsupervised pairwise region analysis method, in the spirit of the PATCH approach. We used the student paintings for validation, dividing them up into 1x1 cm2 (200 x 200 px) patches, giving 180 patches per painting. We then took every possible pair of paintings, to determine whether each pair could be distinguished (marked as “different artist”) or not (marked as “same artist”). For a given pair of paintings (360 patches total) we ran each patch through the following unsupervised workflow:
Feature extraction: each patch was converted to a feature vector using one of two feature extractors: i) a 2D Fourier transform, which created a Fourier domain 200x200 patch with the origin in the center. We took the absolute values of the upper half of the domain (the bottom was identical by symmetry), leaving out the 13 central columns. These columns correspond to the largest wavelengths (comparable to the patch size), and deemphasizing them puts the focus on brushstroke details at smaller wavelengths. We flattened the absolute values into a feature vector of size 17578. Other parameter choices gave comparable or worse final performance; ii) a feature extractor based on the VGG16 convolutional neural network (CNN) that had been trained to classify height data in our earlier student painting controlled experiment 2. The bottom layers of the network were cut out, and the output was pooled and flattened to yield a feature vector of size 512. These two types of feature extractors are labeled as “Fourier” or “CNN” respectively in Table S2.
Dimension reduction: to allow eventual clustering of the feature vectors, we used the UMAP dimension reduction algorithm 46 with default parameters, to map the full feature vectors onto -dimensional vectors. Three different choices of were tested: , , and .
Clustering: Finally, we used the flat clustering module of HDBSCAN 47 (minimum cluster size = 25, cluster selection method = ‘leaf’) to force the algorithm to group the -dimensional vectors into two clusters.
If there was a statistically significant difference in the numbers of patches assigned to each cluster among the two paintings (chi-squared test with p 0.05), then the pair of paintings were marked as different artist. For example, if the labeled patches in one painting were mostly assigned to cluster 1, and the labeled patches in the other painting to cluster 2, this would be taken as evidence that different artists were involved. Otherwise (i.e. if cluster assignments were fairly evenly distributed among the two paintings) the pair was marked as same artist. The algorithm occasionally failed to find two clusters at all, which we also treated as evidence that the pair was made by the same artist. The resulting classification metrics, regardless of feature extractor or dimension , were quite similar to those of the surface roughness approach described above. Changing HDBSCAN parameters like minimum cluster size and the cluster selection method did not substantially alter the performance, and in all cases the unsupervised approach fared worse than the PATCH method.
Supervised UMAP (sUMAP) methods
The final set of methods we tested were supervised variants of the UMAP procedure above. These take advantage of the fact that UMAP can be nudged to provide more distinct separation of clusters by giving it class labels. The approach is the same as earlier, with three differences: 1) 70% of the patches in each of the two paintings are randomly selected to be training data. These patches are given distinct labels by painting (i.e. painting 1 versus painting 2), which are input to the UMAP algorithm along with the patches. 2) The remaining 30% of the patches are used to test whether the approach can successfully distinguish paintings by putting each painting in a distinct cluster. As before, if a statistically significant difference exists in the number of test patches assigned to each cluster among the two paintings (chi-squared test with ) then the pair is marked as different-artist. 3) Because of the stochastic nature of training / test split, the analysis was repeated three independent times, and where there was disagreement we used the majority outcome of the three runs to decide on same-artist or different-artist for each pair.
Both Fourier and CNN feature extractors were tested with this approach, as well as different output dimensions . The supervised component did improve the results, and as seen in Table S2 same-artist precision and different-artist recall were slightly better than PATCH (by 2.9% and 0.3% respectively). However the same-artist recall was relatively poor (0.652 for this method versus 0.87 for PATCH). Hence overall PATCH was still superior in classification, as evidenced by its higher F1 scores for both same-artist and different-artist pairs. Additionally, the computational complexity of this approach (unlike the simpler surface roughness and unsupervised UMAP methods) was comparable to implementing PATCH, so that there were no substantial performance benefits.
Appendix D: Analysis of PATCH-generated Communities
Further analysis of the community structure in the PATCH generated networks can provide additional insight into the community cohesiveness and distinction beyond the modularity index. These properties are quantified by the average community degree, broadly defined as the weighted and normalized sum of all edges between nodes within two community distinctions. This analysis can provide additional insights into the cohesiveness of artistic practice within a single community, and the relative similarity of practice between multiple communities.
The internal degree () is defined as the weighted sum of all edges () between nodes in the same community (), normalized by the theoretical maximum number of edges for available nodes:
The internal degree can be taken as a quantified metric of the intra-community cohesiveness, related to the variability of textures within an identified community. The external degree () between two communities is similarly defined as the weighted sum of all edges between nodes in distinct communities (, ), normalized by the theoretical maximum number of edges for and available nodes in the respective communities:
The external degree can be taken as a quantified metric for the similarity between two separate communities, related to the distinctiveness of textures between them. Note that due to the prior identification of distinct communities, the external degree is expected to be much lower than the internal. Finally, the total external degree for a community () is defined as the external degree between all nodes internal to () and external to () the community, again normalized by the theoretical maximum number of edges:
The total external degree is a quantified metric describing the relative isolation of a single community from the remainder of the network, related to the overall distinction of textures from the whole. In principle, both the pairwise and total degree metrics are useful in PATCH analysis. For instance, a community resulting from extensive conservation work could engender a low total external degree due to the difference in tools and artist purpose, while multiple communities from similar sources, such as multiple students of the same master, could show a similar high external degree describing similarity of practice. The internal degree could describe, for instance, the region-to-region consistency of a single artist, or the conformity of practice by multiple similar artist sources.
Figure S2 shows the computed internal and external network degrees for the PATCH analysis of the two El Greco paintings. Figure S2 A and C show a pairwise comparison between each of the identified communities in each painting, and Figure S2 B and D show the internal degree and total external degree for each community.
In The Baptism, the external degrees between the first three communities are significantly stronger than those involving the fourth community, with the average external degree over four times larger. This indicates a significant isolation of the fourth community (colored orange in Figure 3, and represented by orange bars in Figure S2 B). The strongest external degree is between communities 1 (yellow) and 3 (green) and the weakest is between 3 and 4, which has no connecting edges. Overall, the 3rd community has a slightly greater connection with the remainder of the network, indicating that the height map is the most similar to all other areas. The internal degree is similar for all communities (%), indicating that no community stands out as having a larger disparity in topographical signal. The external degree analysis is largely redundant in the case of Christ on the Cross, since only two communities are proposed. However, it should be noted that the external degree between the two communities (Christ) is significantly larger than for any other two communities in The Baptism except the connection between communities 1 and 3 (Baptism). This indicates that, on average, more similarities can be found between the two communities in Christ on the Cross than those proposed in The Baptism. This is reflected in the low modularity value. Additionally, the internal degree in each community is similar, which can indicate a similar intra-community stylistic distribution, and supports the two-community model.
Appendix E: Brush Construction in the Early Modern Period
In El Greco’s time, brushes were made with either bristles (which are coarse) or hair (which is fine and soft). Each served distinct technical purposes and often came in material-specific shapes 39. Variation could arise from using the hair/bristles of different species in the same type of brush. Key sources were hog bristles (especially white hog) and “minever” (gray squirrel tail fur) 41 , though dog and mongoose often replaced minever due to its rarity 39; 40. Hog’s bristles likewise may have been replaced with ox 48. Such changes could affect paint application. For example, hog’s bristles are naturally flagged (the ends of the hairs are split), allowing them to pick up more paint 49. Bristle substitution could decrease brush capacity. Further experimentation to understand the effect of materials on PATCH analysis is needed.
Appendix F: Additional Figures
Appendix G: Edge Pruning
The PATCH method is concerned with “sameness,” meaning we need a way to remove spurious edges from the network. To accomplish this, we employ an edge pruning process based on uniqueness. We assume that regions with a high degree (number of connections) have more common features and nodes with a low degree have more unique features. Therefore, a connection between regions with more unique features is more likely to be correct. We operationalize this assumption by assigning each edge a score based on the average uniqueness of its nodes, where uniqueness is defined as the percentage of nodes a node does not connect with: where max = the maximum possible degree (i.e., number of connections) and degree = the degree of one of the two nodes in question. There were two erroneous same-artist pairs in our experimental dataset (9% of edges in the network), and these edges had the lowest uniqueness scores. Thus, we trim the 9% of edges (or more if there are ties) in a given network with the lowest uniqueness scores.